The function gleft(xright)=fracx−−35x−7is not defined for x=___________________.
a.7
b.\\(frac{5}{7}\\)
c.5
d.\\(frac{7}{5}\\)
1
Expert's answer
2020-05-19T19:07:33-0400

The function "f(x)=\dfrac{x-35}{x-7}."


"x-7\not=0=>x\not=7"

The function "f(x)=\dfrac{x-35}{x-7}" is not defined for "x=7."

a. 7